the perimeter of the rectangle is 60 inches the length is 8 inches greater than width what is width and length
Answer by Cromlix(4381)

Width = x
Length = 8 + x
Perimeter = 2*width + 2*length
60 = 2(x) + 2(8 + x)
60 = 2x + 16 + 2x
4x = 60 - 16
4x = 44
x = 11
Width = 11 inches
Length = 19 inches
Hope this helps.
:-)